Could you please help me with this question:Christopher Columbus just spent 20 pieces of silver at the Mariner's Mall. He bought a hat, a cape, and a spyglass. The spyglass cost 9 more pieces of silver than the cape. The spyglass and the cape together cost 16 more pieces of silver than the
hat. How much did Christopher Columbus pay for his hat?

Answer by mananth(12270)

Let cape cost x
spy glass , x+9
Hat = (2x+9-16)=2x-7
x+x+9+2x-7=20
4x+2=20
4x=18
x=4.5--------- cape
spy glass ---- 9+4.5 = 13.5
Hat = 2